A manufacturing process has 100 customer orders to fill. Each order requires one component part that is purchased from a supplier. However, typically, 2% of the components are identified as defective, and the components can be assumed to be independent.
a. If the manufacturer stocks 100 components, what is the probability that the 100 orders can be filled without reordering components?
b. If the manufacturer stocks 102 components, what is the probability that the 100 orders can be filled without reordering components?
c. If the manufacturer stocks 105 components, what is the probability that the 100 orders can be filled without reordering components?

Answer by stanbon(75887)

A manufacturing process has 100 customer orders to fill. Each order requires one component part that is purchased from a supplier. However, typically, 2% of the components are identified as defective, and the components can be assumed to be independent.
a. If the manufacturer stocks 100 components, what is the probability that the 100 orders can be filled without reordering components?
----
Ans: 0.98^100 = 0.1326
================================
b. If the manufacturer stocks 102 components, what is the probability that the 100 orders can be filled without reordering components?
Ans: 102C100*0.98^100*0.02*2 = 0.2732
================================
c. If the manufacturer stocks 105 components, what is the probability that the 100 orders can be filled without reordering components?
Ans: 105C100*0.98^100*0.02^5 = 0.0410
================================
Cheers,
Stan H.
=============